the half life of chromium-56 is 28 days, if the sample contained 150 grams, how much was present 168 days ago


1
Expert's answer
2021-09-16T14:19:19-0400

Let’s use the formula for radioactive decay:


"N=N_0e^{-\dfrac{0.693}{T_{1/2}}t},""N=150\ g\cdot e^{-\dfrac{0.693}{28\ days}\cdot168\ days}=2.35\ g."